Hi!
Help me please! I don't understand how determine an indefinite multiple integral. For instanse, \iint 1 dxdy = ?
1
Expert's answer
2020-03-03T18:07:40-0500

"\iint1dxdy=F(x,y)" ,where "F(x,y)=y(x+c_1)+c_2" , where "c_1=const,c_2=const" , because "\frac{d^2F(x,y)}{dxdy}=1"